### v1

# Model Provenance Testing for Large Language Models

Abstract
Large language models are increasingly customized through fine-tuning and other adaptations, creating challenges in enforcing licensing terms and managing downstream impacts. Tracking model origins is crucial both for protecting intellectual property and for identifying derived models when biases or vulnerabilities are discovered in foundation models. We address this challenge by developing a framework for testing model provenance: Whether one model is derived from another. Our approach is based on the key observation that real-world model derivations preserve significant similarities in model outputs that can be detected through statistical analysis. Using only black-box access to models, we employ multiple hypothesis testing to compare model similarities against a baseline established by unrelated models. On two comprehensive real-world benchmarks spanning models from 30M to 4B parameters and comprising over $600$ models, our tester achieves $90-95$% precision and $80-90$% recall in identifying derived models. These results demonstrate the viability of systematic provenance verification in production environments even when only API access is available.
1 Introduction
Platforms such as Amazon SageMaker and Hugging Face have enabled wide scale distribution of ML models, most notably large language models (LLMs) [40, 50, 22]. Certain models, called foundational models, require extensive computational resources and datasets to train. For instance, it was reported in [1] that Meta used two $24,000$ GPU clusters to train Llama 3. But foundational models are relatively cheap to fine-tune or customize for downstream applications [25, 51]. We are thus seeing a proliferation of fine-tuned models.
The increase of publicly available foundation models and datasets, however, has also triggered concerns over unauthorized use of intellectual property (IP) and concerns of compromised datasets [10] or models [20]. These issues are present not just in open-source ecosystems, but also for proprietary models that are hidden behind APIs [4, 33]. For instance, concerns about model stealing attacks wherein one can extract the model parameters even for production-level models are on the rise [11, 47, 34]. Similarly, there is a growing concern that proprietary models may contain backdoors or vulnerabilities, making them susceptible to jailbreaking [20, 3, 56]. Despite best efforts to create a safe environment for the development of (public and commercial) foundation models, there have already been instances of reported misuse [38, 18, 46].
This landscape highlights the growing need for model provenance testing. The problem of model provenance testing is as follows: Determine whether a target model has been derived from a foundational model by lighter customization such as fine-tuning. This problem has applications in tracking reuse of models not just in open marketplaces but also across product teams in large organizations. Current policy regulations such as GDPR [14] and Artificial Intelligence (AI) Safety Act [13] require security and privacy compliance for all AI-enabled apps [16]. When a security or privacy audit finds a problem with a foundational model, it becomes important to identify which other models in use by the organization may be derived from the problematic one and take appropriate remedial actions (e.g. revoke, retrain, or fortify) to mitigate the risk of further non-compliant use. Model provenance tracking is often useful after the fine-tuned model has been deployed, and when authentic ground truth is unavailable or unreliable.
One challenging aspect of designing a model provenance tester is achieving high accuracy. There is a cost associated with a provenance verdict. For instance, as a result of provenance tracking, a company may initiate legal action or investigation. For use cases within the same organization, developers might have to revoke the use of an existing model and even retrain from a clean parent model. False positives, i.e., the deployed LLM is wrongly flagged as a derivation of a problematic LLM, thus entail a downstream cost. At the same time, false negatives, i.e., not being able to flag that the LLM is customized from a problematic parent, also increase the risk of non-compliance. Therefore, we want a principled way to decide provenance and to make accuracy trade-offs.
Another challenge is that a practical provenance tool needs to have minimal assumptions to be readily usable in many post-deployment settings. We focus on techniques that do not change typical training and data pipelines, and can be integrated for current state-of-the-art LLMs. The tester is expected to only have black-box query access to the models and has no additional information, such as the training dataset, test set, or the algorithm used for training. We are not aware of any prior work addressing the question of model provenance testing systematically and in such practical setups.
In this paper, we design the first practical model provenance tester for LLMs that requires only query access. Our proposed techniques stem from a key empirical observation: The output distribution of fine-tuned models is often close to that of their parent model. This distance between a model and its true parent is typically smaller than that between the model and another unrelated models, making it possible to reliably trace back a derived model to the original parent. In order to keep assumptions to a minimum, we propose to employ the principled framework of statistical hypothesis testing. Specifically, we use black-box sampling and estimation to determine whether the distribution captured by the given model is close to that of the candidate parent. Such estimation can provide formal statistical significance measures, which can be used to check for the null hypothesis, i.e., the customized LLM is not close to the given parent model.
We conduct an extensive empirical evaluation across two comprehensive benchmarks comprising over $600$ models from Hugging Face, ranging from $30$M to $4$B model parameters. Our tester achieves $90-95$% precision and $80-90$% recall in detecting model provenance relationships, even with a limited number of queries. Therefore, while being simple, we find that our proposed method achieves excellent accuracy.
Contributions. We present the problem of model provenance testing. Our work initiates the study of such testers in the context of customized LLMs, and keeps assumptions minimal, i.e., having only black-box query access to models. We show that in the existing landscape of open-source LLM models, it is practical to fairly accurately determine provenance.
2 Overview
Model provenance testing has many motivating applications, but we present one representative scenario for concreteness.
2.1 Motivating Scenario
Pretraining large language models (LLMs) involves significant investment, requiring substantial computational resources costing millions of dollars in infrastructure and thousands of GPU hours. When Company A releases a pretrained LLM (called LLM-$f$), it employs specific licensing terms crucial for protecting this investment, maintaining competitive advantage, and controlling the model’s usage [33, 4]. These licenses typically include restrictions on commercial usage, model modification, or redistribution, and may incorporate provisions for monitoring downstream usage and revenue sharing.
Startup B might download LLM-$f$, perform only fine-tuning, but claim to have pretrained their model (LLM-$g$) from scratch, thereby circumventing licensing requirements and misrepresenting company A’s work. In such cases, we want to be able to determine if LLM-$g$ is derived through fine-tuning LLM-$f$ and resolve the model provenance problem. Given the significant economic incentive to fine-tune existing models rather than pretrain from scratch, verifying model provenance becomes essential for enforcing licensing terms and protecting intellectual property rights. Aside from fine-tuning, there are a few other inexpensive techniques (mixture-of-experts, prompt engineering) that Startup B could employ to derive LLM-$g$ from LLM-$f$ while still claiming independent development. All these approaches require orders of magnitude less computational resources than pretraining, making them particularly attractive for circumventing licensing restrictions. Thus, the model provenance problem extends beyond detecting fine-tuning to identifying any form of model derivation that might violate licensing terms. We have witnessed such a case recently when the pretrained model Llama 3.1 was reportedly used for military operations, violating the license agreement [38].
Startup B can release LLM-$g$ either by making its weights publicly available (e.g., on Hugging Face) or by offering it as an online service through an API (e.g., ChatGPT or PaLM-2). The API-like interface is more general than if weights were also available to a tester trying to determine provenance. In this setup, the tester must rely solely on the model’s responses to input queries. This distinction significantly impacts the techniques available for detecting whether LLM-$g$ is derived from LLM-$f$.
We consider a testing framework with minimal assumptions only, where the tester can query LLM-$g$ on arbitrary prompts (through the API) and get responses. The tester has no access to the training datasets, hyperparameters used by either company, or any information about potential modifications performed by Startup B. This mirrors real-world conditions where companies do not always disclose their training procedures, data sources, or modification techniques, making the provenance testing problem both practical and challenging.
2.2 Our Approach
Figure 1: Our model provenance tester that decides if model $g$ is derived from model $f$.
Our approach to testing model provenance is based on a key observation: fine-tuning and other model derivation techniques typically result in only limited changes to the original model, as they primarily adapt the model for new tasks. After fine-tuning, the derived model LLM-$g$ may remain similar to its parent model LLM-$f$, as the process focuses on refining specific capabilities rather than creating fundamental changes to the model distribution.
In our setting, the tester can only interact with models through their API interfaces, submitting prompts and analyzing the generated outputs. We thus compute similarity of models by analyzing the behavioral patterns between LLM-$f$ and LLM-$g$ through their responses. For efficiency and simplicity, we focus on comparing next-token predictions on random prompts rather than analyzing entire sequences. While extending the analysis to longer sequences ($n$-grams) might provide additional signals, our empirical evidence suggests that the next-token distribution already captures significant information about a model’s decision-making patterns and internal representations.
Our method tests a number of random diverse prompts, looking for evidence that LLM-$g$’s output distributions consistently align more closely with LLM-$f$ than would be expected from a model that was not derived from LLM-$f$. However, there is no fixed threshold for what constitutes “close alignment”. We cannot simply say that matching on 10% or 20% of outputs indicates provenance, as the degree of similarity can vary significantly depending on the intensity of fine-tuning, how different the downstream task is and other modifications performed. Even models from which LLM-$g$ was not derived may occasionally produce matching outputs, and we need to understand this rate of coincidental alignment. We thus need more rigorous and universal approach, that is based not only on similarity between models, but on finding a level of similarity that is unusually higher than coincidence.
To establish a meaningful baseline, i.e., to characterize the expected level of coincidental agreement between unrelated models, we take several additional control models and analyze how closely their outputs align with LLM-$g$. The selection of control models is crucial for establishing reliable baselines for coincidental agreement. These control models are selected to represent a diverse range of output distributions, thus allowing to better estimate the average levels of similarity that might occur between unrelated models. If the similarity between LLM-$f$ and LLM-$g$ is substantially higher than the similarity observed with the control models, it provides compelling evidence that LLM-$g$ was indeed derived from LLM-$f$, as their behavioral patterns exhibit a level of consistency that cannot be easily attributed to random chance. Conversely, if the similarity falls within the typical range observed among the control models, it may indicate that the observed alignment is not sufficient to conclusively determine provenance.
The actual test for excess similarity between LLM-$f$ and LLM-$g$ is performed using multiple hypothesis testing. Each hypothesis is formulated to show that similarity of LLM-$g$ to LLM-$f$ is greater than the similarity to any control models. When all these hypothesis tests yield statistically significant results, we obtain strong evidence that the similarity between these models is systematically higher than what would be expected by chance. This statistical framework offers several key advantages. First, it handles varying degrees of fine-tuning adaptively, without requiring assumptions about the expected similarity between LLM-$g$ and LLM-$f$. The test naturally identifies any significant bias in matching rates compared to the best non-parent baseline model. Second, rather than relying on arbitrary thresholds, it provides a principled way to quantify confidence in provenance determination through $p$-values, i.e., observed bias in matching rates, and confidence intervals. Third, the approach is computationally efficient, requiring only a modest number of test prompts to achieve statistical significance.
2.3 Advantages and Effectiveness
Our approach stands out in its reliance on minimal assumptions, particularly the requirement of only black-box query access to the models in question. By operating without access to proprietary model weights, training data, or hyperparameters, we mirror real-world conditions where such internal details are typically inaccessible. This makes our method highly practical and broadly applicable, as it can be employed in scenarios where models are available exclusively through API endpoints with restricted internal information.
A key strength of our method is its statistical foundation. By employing multiple hypothesis testing, we objectively assess whether the observed similarities between LLM-$g$ and LLM-$f$ are significant beyond what could be attributed to random chance. This framework allows us to compute precise $p$-values and confidence intervals, providing quantifiable evidence to support our conclusions about model similarities. Importantly, our approach avoids reliance on arbitrary thresholds and naturally adapts to varying degrees of similarity.
From a practical standpoint, our method is both efficient and scalable. Focusing on next-token predictions over a diverse set of random prompts minimizes the computational overhead and the number of queries required. This efficiency is particularly valuable when dealing with large models and the inherent limitations of API access, such as query costs and rate limits. The simplicity of our approach allows for straightforward implementation without the need for specialized hardware or extensive computational resources.
Regarding effectiveness, our method has demonstrated high accuracy in detecting model provenance. As we will detail in Section 4, our approach consistently identifies provenance pairs with over $90$% precision and $80$% recall when tested on large set of pairs. These results highlight the method’s capability to reliably determine provenance, even when the derived model has been significantly fine-tuned or altered. This effectiveness underscores its practical utility in real-world applications where accurate provenance detection is critical.
2.4 Challenges
While our approach offers a robust and practical solution to the model provenance problem, it is important to acknowledge its limitations and the challenges inherent in this domain. Understanding these factors is crucial for interpreting the results of our method and guiding future enhancements.
First, a primary challenge is dealing with cases where the derived model LLM-$g$ has undergone extensive modifications that significantly alter its behavior relative to the parent model LLM-$f$. Such modifications might include heavy fine-tuning on large and diverse datasets, architectural changes, or techniques designed to intentionally mask the model’s origins. In these scenarios, the residual similarities between LLM-$g$ and LLM-$f$ may diminish to the point where our statistical tests lose sensitivity, potentially leading to false negatives.
Second, our method depends on the selection of control models used for baseline comparisons. The effectiveness of our statistical tests hinges on having appropriate control models that are sufficiently diverse and representative of unrelated models. If the control models are inadvertently dissimilar to LLM-$f$, this could affect the baseline similarity levels and impact the test’s reliability. Careful consideration is therefore required in choosing control models to ensure they provide a meaningful contrast in the hypothesis testing framework.
Another challenge lies in the inherent variability of language models and the possibility of coincidental similarities. Language models trained on similar datasets might exhibit overlapping behaviors, even if one is not derived from the other. This could potentially lead to false positives, where our method incorrectly identifies a provenance relationship due to shared patterns that are common in the domain rather than indicative of direct derivation. Addressing this requires careful interpretation of the results and may need an additional analysis to rule out such coincidental similarities.
In Section 4 we will experimentally determine the impact of some of these assumptions on the tester’s effectiveness.
3 Main Approach
In this section we provide a more formal introduction of the problem and present the main model provenance tester.
3.1 Problem Setup
A large language model (LLM) takes as input a sequence of tokens $x_{<i}=(x_{1},\ldots,x_{i-1})$ from a fixed set of tokens $\mathcal{V}$ and outputs the next token $x_{i}\in\mathcal{V}$, i.e., $f_{\theta}(x_{<i})=x_{i}$. Formally, LLM is denoted with $f_{\theta}$, defined by its architecture $f$ and its model parameters $\theta$. Herein, it is sufficient for us to consider the LLM as a mapping $f_{\theta}:\mathcal{V}^{I}\rightarrow\mathcal{V}$, where $I$ is an arbitrary sequence length. Conditioned on an input sequence $x_{<i}$, the $f_{\theta}$ induces a probability distribution on the whole vocabulary set $\mathcal{V}$ so each token is assigned a probability score. In the sequel, we assume $f_{\theta}$ chooses the token with the highest probability, so-called temperature $t=0$ case. Note, when $f_{\theta}$ uses higher temperatures $t>0$, the most likely token can still be empirically determined by repeatedly sampling from the induced distribution and selecting the most frequently occurring token.
A model $f_{\theta}$ can be customized through various techniques to create a new model. The most common approach is fine-tuning, where the original parameters $\theta$ are updated using gradient descent on a new dataset to obtain $\theta^{\prime}$. Other customization techniques such as mixture-of-experts can involve creating a model $g_{\theta^{\prime}}$ with different architecture that reuses parts of the architecture of the initial model $f_{\theta}$. Thus, in the sequel, when addressing customizations and models, we omit parameter notations, and denote the initial (parent) model as $f$, and the derived (child) model as $g$, and say that $(f,g)$ constitutes a model provenance pair. Importantly, these customizations typically require orders of magnitude less computational resources than training from scratch. Under our setup, we have only query access to the models - we can provide input sequences $x_{<i}=(x_{1},\ldots,x_{i-1})$ and observe their corresponding output tokens $x_{i}$. This reflects real-world scenarios where models are accessed through APIs, while keeping internal model details private. The model provenance problem consists thus in detecting if $g$ was derived from $f$, given only query access to models.
Threat Model.
Our proposed methods are primarily evaluated for a non-adaptive adversary. The model customizer is not aware of the strategy of the model provenance tester. They are not aware of the knowledge the tester has, i.e., the control models, the set of randomly sampled inputs, or the parameters of the statistical test. On the other hand, one may consider an adaptive adversary, i.e., customizers that may employ techniques specifically designed to evade provenance detection, especially based on the knowledge of the provenance tester. For instance, they might introduce noise or deliberate alterations to the output distributions to reduce detectable similarities with the parent model. Such adaptive obfuscation pose a significant challenge, as they can diminish the effectiveness of our statistical tests. Developing methods to counteract such adaptive adversarial strategies is beyond the scope of this work.
3.2 Model Provenance Tester
We formulate the model provenance question as the problem of checking whether two models are similar. This approach is based on the observation that models derived from one another tend to exhibit higher similarity compared to unrelated models. We now detail how the similarity check is implemented.
Since the tester has only query access to the models, the only data it can collect is from providing inputs (called prompts) and analyzing the corresponding output tokens. Furthermore, due to our minimal assumptions and lack of information about the training datasets, the tester queries the models on randomly chosen short sequences designed to avoid predictable (low-entropy) continuations. For example, the tester avoids prompts like “To be or not to” since the next token is highly likely to be “be” a common continuation that many models would predict. [^1] Tokens with low entropy are those for which the next token can be easily predicted by any model, making them unsuitable for distinguishing between models. We further denote such prompt space with $\Omega$ and sample from it uniformly at random.
The tester independently samples from $\Omega$ a set of $T$ such prompts, queries each model on the same set of prompts, and then compares their output tokens pairwise. For each prompt, we compare only the first output token generated by each model; however, in principle, $n$-grams (sequences of $n$ consecutive tokens) could also be considered. The similarity between two models is then calculated as the proportion $\mu$ of prompts on which the models produce the same output token:
$\mu=\frac{1}{T}\sum_{j=1}^{T}\mathds{1}(f(x_{<i}^{j})=g(x_{<i}^{j}))$
However, the ratio $\mu$ by itself may not be sufficient to determine whether $f$ and $g$ are truly similar. We need to understand if the similarity $\mu$ between $g$ and $f$ is higher than what we would expect if $g$ were unrelated to $f$. To make this assessment, we need to establish what level of similarity we should expect between $g$ and models that are unrelated to it but share key characteristics with $f$. We do this by introducing a set of control models $C=\{c_{1},\ldots,c_{m}\}$ chosen specifically for $f$. These control models are selected to match $f$’s domain, purpose or capabilities as closely as possible, while being clearly not ancestors of $g$. They are selected from publicly available models that share similar capabilities with $f$ but have been independently developed using different training datasets, architectures, or approaches. Such suitable control models are readily available in existing model repositories with sufficient variability in their development approaches. For instance, if $f$ is a French-language model for generating Python code, appropriate control models would include other French language models, Python code generation models, or ideally other French Python code generation models. On the other hand, if $f$ is a more general LLM such as Llama-3.2-1B, then the control models would include other general LLMs, such as Llama-3.2-3B, GPT2-large and Qwen2-1.5B. We do not have knowledge and influence beyond the domain or capabilities (e.g., “general” vs. “’French”). The number $m$ of control models depends on availability; generally, more control models from $f$’s domain lead to better baseline estimation. We observe empirically that there is enough diversity in existing pre-trained models such that constructing the control set does not require significant computational effort (e.g., additional training) or specialized machine learning expertise. Moreover, if the domain of the parent model $f$ is unknown, one can deduce its domain by manually analyzing the outputs of $f$ on some sample prompts. If such manual analysis is not feasible, we can just consider a larger and more diverse set of LLMs as the control set. For each control model $c_{i}$, we compute its similarity ratio $\mu_{i}$ with $g$ using the same process as before. These $\mu_{i}$ values form our baseline - they tell us how much $g$ typically agrees with models that operate in the same space as $f$ but are definitely not its ancestors. If the similarity $\mu$ between $g$ and $f$ significantly exceeds these baseline similarities $\mu_{i}$, this suggests a provenance relationship.
The final step of the tester is to verify that the similarity ratio $\mu$ between $f$ and $g$ exceeds all similarity ratios $\mu_{i}$. However, we want to ensure this is not merely due to random chance, but rather reflects a true difference. To establish such theoretical guarantees, we employ multiple hypothesis testing. More precisely, for each control model $c_{i}$, we formulate the following hypothesis test $H^{i}$:
$H_{0}^{i}:$ $\quad\mu\leq\mu_{i},$
$H_{1}^{i}:$ $\quad\mu>\mu_{i},$
where $H_{0}^{i}$ is the null hypothesis that the similarity between $f$ and $g$ is less than or equal to the similarity between $c_{i}$ and $g$, and $H_{1}^{i}$ is the alternative hypothesis that the similarity between $f$ and $g$ is greater. To test each of the hypothesis $H^{i}$, we employ a z-test, which is a standard statistical test well-suited for comparing proportions like our similarity ratios when working with large samples. The z-test helps us determine if the observed difference between the two proportions ($\mu$ and $\mu_{i}$) is large enough to be statistically significant, or if it could have occurred by chance. The z-test produces a $p$-value, which represents the probability of observing such a difference in proportions if there were truly no difference between the models (i.e., if the null hypothesis were true). A small $p$-value (typically less than a predetermined significance level $\alpha$) suggests that the observed difference is unlikely to have occurred by chance, indicating that the similarity $\mu$ between $f$ and $g$ is indeed significantly larger than the similarity $\mu_{i}$ between $g$ and the control model $c_{i}$. We want all hypothesis tests to yield small $p$-values, indicating that the similarity $\mu$ is significantly higher than every baseline similarity $\mu_{i}$.
The significance level $\alpha$ represents our tolerance for incorrectly concluding that a difference is significant when it actually occurred by chance (a false positive). This threshold is set to a commonly used default value of $0.05$, meaning we accept a $5\%$ risk of claiming a significant difference where none truly exists. In our context, this would mean wrongly concluding that $\mu$ is significantly higher than some $\mu_{i}$ when the observed difference is merely due to chance.
When conducting multiple hypothesis tests simultaneously, we want to maintain this same overall risk level of $\alpha$, regardless of how many tests we perform. However, running multiple tests increases our chances of obtaining at least one false positive across all tests (known as the family-wise error rate FWER). To control this cumulative risk, we employ the Holm-Bonferroni method [19], which adjusts the significance thresholds $\alpha_{k}$ for individual tests $H^{k}$ to ensure the overall false positive rate remains at or below our desired level of $\alpha$. The Holm-Bonferroni procedure works as follows. We first sort the individual $p$-values in ascending order: $p_{(1)}\leq p_{(2)}\leq\dots\leq p_{(m)}$. We then compare each p-value $p_{(k)}$ to its adjusted significance level $\alpha_{k}=\alpha/(m-k+1)$. Starting with the smallest $p$-value, we sequentially test each hypothesis. If a $p$-value $p_{(k)}$ is less than or equal to its corresponding $\alpha_{k}$, we reject the null hypothesis $H_{0}^{(k)}$ and proceed to the next one. This process continues until we encounter a $p$-value that exceeds its adjusted significance level, at which point we fail to reject the remaining null hypotheses. To conclude that $\mu$ is significantly greater than all $\mu_{i}$, we need to reject all null hypotheses $H_{0}^{i}$. By carefully controlling the FWER and ensuring that all null hypotheses are rejected through this method, we provide strong statistical evidence that the similarity between $f$ and $g$ is higher than that between $g$ and any control model $c_{i}$. This supports the assertion that $(f,g)$ has significantly higher similarity and thus constitutes a model provenance pair.
Algorithm 1 Model Provenance Tester for Pair $(f,g)$
Pair $(f,g)$, set of control models $C=\{c_{1},\ldots,c_{m}\}$, prompt space $\Omega$, number of prompts $T$, significance parameter $\alpha$, statistical test ZTest.
$x_{1},\ldots,x_{T}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\Omega$triangleright Sample T prompts
$\mu\leftarrow\frac{1}{T}\sum_{j=1}^{T}\mathds{1}(f(x_{j})=g(x_{j}))$triangleright Calc sim of f and g
for $i\leftarrow 1$ to $m$ do
     $\mu_{i}\leftarrow\frac{1}{T}\sum_{j=1}^{T}\mathds{1}(c_{i}(x_{j})=g(x_{j}))$triangleright Calc sim of $c_{i}$ and $g$
     $p_{i}\leftarrow\text{ZTest}(\mu,\mu_{i},T)$triangleright Obtain p-values
end for
$(p_{(1)},\ldots,p_{(m)})\leftarrow\text{Sort}(p_{1},\ldots,p_{m})$triangleright Sort p-values
for $k\leftarrow 1$ to $m$ do
     $\alpha_{k}\leftarrow\alpha/(m-k+1)$triangleright Holm-Bonferroni adjustment
     if $p_{(k)}>\alpha_{k}$ then return False $\triangleright$ Not a provenance pair
     end if
end forreturn True $\triangleright$ Is a provenance pair
A pseudo-code of the tester is given in Algorithm 1. In summary, when the procedure returns True, it ensures that the total family-wise error rate (FWER) is controlled at the significance level of $\alpha=0.05$ or lower. This means we can confidently state that the similarity between $f$ and $g$ is significantly higher than between $g$ and any control model, supporting the existence of a provenance relationship. Conversely, when the procedure returns False, it indicates that we could not establish this higher similarity with the desired level of statistical significance. This may occur either because there is genuinely no significant similarity indicative of provenance or because the test lacked sufficient power under the given parameters (e.g., sample size or number of prompts) to detect it.
Algorithm 2 Provenance Tester for $g$ Given a Candidate Parent Set
Model $g$, candidate set $F=\{f_{1},\ldots,f_{s}\}$, set of control models $C=\{c_{1},\ldots,c_{m}\}$, prompt space $\Omega$, number of prompts $T$, significance parameter $\alpha$, statistical test ZTest.
$x_{1},\ldots,x_{T}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\Omega$triangleright Sample T prompts
for $i\leftarrow 1$ to $s$ do
     $\mu_{i}\leftarrow\frac{1}{T}\sum_{j=1}^{T}\mathds{1}(f_{i}(x_{j})=g(x_{j}))$triangleright Calc sim of candidates
end for
for $i\leftarrow 1$ to $m$ do
     $\mu^{\prime}_{i}\leftarrow\frac{1}{T}\sum_{j=1}^{T}\mathds{1}(c_{i}(x_{j})=g(x_{j}))$triangleright Calc sim of controls
end for
$\mathcal{M}\leftarrow\{\mu_{1},\ldots,\mu_{s}\}\cup\{\mu^{\prime}_{1},\ldots,\mu^{\prime}_{m}\}$triangleright Set of all sims
$\mu_{max}\leftarrow\max(\mathcal{M})$triangleright Find highest sim
if $\mu_{max}\notin\{\mu_{1},\ldots,\mu_{s}\}$ then return False $\triangleright$ Highest not from $F$, but from $C$, so cannot be parent
end if
for $\mu^{\prime}\in\mathcal{M}\setminus\{\mu_{max}\}$ do
     $p_{i}\leftarrow\text{ZTest}(\mu_{max},\mu^{\prime},T)$triangleright Compare against other sims
end for
$(p_{(1)},\ldots,p_{(s+m-1)})\leftarrow\text{Sort}(p_{1},\ldots,p_{s+m-1})$
for $k\leftarrow 1$ to $s+m-1$ do
     $\alpha_{k}\leftarrow\alpha/(s+m-k)$triangleright Holm-Bonferroni adjustment
     if $p_{(k)}>\alpha_{k}$ then return False
     end if
end forreturn $(\textsc{True},\operatorname*{arg\,max}_{i\in[s]}\mu_{i})$triangleright Return parent
Extended Model Provenance Test.
The above provenance problem described thus far assumes we have one candidate parent $f$ that we want to test our child model $g$ against. We now extend it to the general case when there is a set of parents, which we refer to as the extended model provenance problem (with unspecified parent). In this problem, given only query access to models, the goal is to determine whether a model $g$ is derived from some model from the set ${f_{1},\ldots,f_{s}}$ of candidate parent models.
While running the basic tester $s$ times (once for each provenance pair $(f_{i},g)$) would solve the extended parent problem, this approach besides requiring more effort, also would require additional correction for multiple testing to maintain the same level of confidence. The probability of false positives would grow with the number of candidate parents $s$ unless appropriate adjustments (such as Holm-Bonferroni correction) are made to the significance level. We thus consider improved tester given in Algorithm 2. Our tester avoids this issue by conducting a single set of hypothesis tests after identifying the most similar candidate. It works as follows. First it finds the most similar model to the given model $g$ among all the control models $C$ and candidate parents $F$. If that model is a control model, the algorithm terminates with False. Otherwise, it goes on to test whether the FWER of this model is overall below $\alpha$, the desired significance level. The test for the latter is the same as in Algorithm 1, except now all the alternate hypotheses (including control and candidates) are in the family tested against. When the algorithm return True, it has the guarantee that the most similar model is one of the candidate models and that the total significance level across all hypotheses meets the threshold $\alpha$.
3.3 Understanding Sources of Error
Our provenance tester can make two types of errors: false positives, where it incorrectly identifies a provenance relationship between independently developed models, and false negatives, where it fails to detect an actual derivation relationship between models.
These errors arise in part from our statistical hypothesis tests. When performing such tests, we can control false positives by setting a stricter significance level $\alpha$ (requiring stronger evidence before claiming excess similarity), while we can reduce false negatives by increasing the sample size $T$ (more samples provide better power to detect true similarities).
Besides errors introduced by statistical tests, our approach of testing provenance through similarity detection may introduce additional errors. Since we claim provenance only when we detect higher-than-expected similarity between models, we need to examine what this approach implies for our error analysis. This approach relies on two key assumptions that can impact error rates:
Assumption 1: Derivation implies similarity.
We assume that when model $g$ is derived from $f$, they will exhibit above-average similarity in their outputs. This assumption leads to two potential types of errors:
  • False negatives occur when a derived model shows insufficient similarity to its parent. This can happen when a model customizer applies extensive modifications that significantly alter the model’s behavior. While resource constraints typically prevent such extreme modifications (as they would approach the cost of training from scratch), some legitimate derivation relationships may still go undetected.
  • False positives arise when independently developed models exhibit high similarity. This typically happens when models are trained on similar datasets or designed for similar specialized tasks - for instance, two independent medical diagnosis models may produce very similar outputs due to their shared domain constraints.
Assumption 2: Control models establish a valid baseline.
We assume our control models provide a reliable baseline for the similarity we should expect between unrelated models. Poor selection of control models can lead to two types of errors:
  • False positives occur when our control models are too dissimilar from the domain of $f$. For example, using general language models as controls for specialized code generation models sets an artificially low baseline, making normal domain-specific similarities appear significant.
  • False negatives happen when control models are themselves derived from $f$ or trained on very similar data. This establishes an artificially high baseline that masks genuine derivation relationships.
The overall error rates of our tester depend on the combination of errors from both our statistical hypothesis testing and the two core assumptions. While we can provide theoretical guarantees for controlling error rates in hypothesis testing through parameters $\alpha$ and $T$, we cannot derive analytical bounds for errors arising from the assumptions about derivation implying similarity or the validity of the control model baseline. These assumption-based error rates can only be evaluated empirically. However, our extensive experiments in Section 4 demonstrate that these assumptions hold well in practice across a wide range of models, suggesting that our approach of provenance testing to similarity detection is sound for real-world applications with non-adaptive adversaries.
3.4 Reducing Query Complexity
Most of LLMs available currently allow cheap (even free) API access, thus the monetary query cost of running our testers is insignificant. When this is not the case, for example, either when the cost of queries is high (e.g. one query to OpenAI model O1 can cost more than \$1 [32]), or the models have some rate restrictions, one can consider enhancements to our testers from Algorithms 12. Furthermore, there are use cases when query complexity can be reduced without any side effects, thus it makes sense from optimization perspective. Note that our proposed enhancements for query reduction are not meant to preserve the theoretical guarantees of classical hypothesis testing that our previous testers inherit, but they can be useful in setups where query costs are prohibitive.
We can divide the queries used in the tester (see Algorithms 12) into two distinctive types: online queries made to the tested child model $g$, and offline queries made to the parent model $f$ (or models $f_{1},\ldots,f_{s}$) and to the control models $c_{1},\ldots,c_{m}$. We make this distinction for two reasons. First, often offline queries are much cheaper, as the potential parent models (and the control models as we will see in the Section 4) are well established, and available from multiple sources, thus they are usually cheaper or free. Second, in some use cases, we can reuse the offline queries to perform many provenance tests of different $g_{i}$. Thus further we analyze separately these two scenarios.
Reducing Online Complexity. Since our tester is fundamentally based on statistical hypothesis testing, any reduction in query complexity must be compensated by increasing the statistical power of individual queries. Rather than querying model $g$ with $T$ random prompts, we can strategically select a smaller set of $T^{\prime}<T$ prompts that yield comparable statistical power for detecting model provenance [^2] It means in Algorithms 12, instead of random sampling $x_{1},\ldots,x_{T}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\Omega$, the goal is to find set $x_{1},\ldots,x_{T^{\prime}}$ from $x_{1},\ldots,x_{T}$ and $F,C$.. We achieve this through an informed sampling approach: instead of uniform sampling from $\Omega$, we employ rejection sampling with an entropy-based selection criterion. Specifically, to generate each prompt in $T^{\prime}$, we sample $k$ candidate prompts from $\Omega$ and select the one that maximizes the entropy of output tokens across all parent and control models. The selection criterion is dynamically weighted to favor prompts that have stronger discriminative power between similar models. While this approach introduces dependencies between the sampled prompts (so the theoretical guarantees of classical hypothesis testing used in Algorithm 1 and 2 do not carry over), our empirical results in Section 4.3 demonstrate its practical effectiveness. Full details about the approach are given in Appendix E.
Reducing Offline Complexity. In non-adversarial settings where multiple provenance tests are performed against the same parent model $f$, we can trivially reduce the offline complexity by reusing the same set of offline queries across all tests. A concrete example of such scenarios arises when issues are discovered in a pre-trained LLM, such as problematic training data or generation of harmful content. A recent example is the lawsuit against the pre-trained model LLama for using copyrighted data in its training set [7]. Since various teams and organizations may have fine-tuned their applications using this model, but precise provenance information is not readily available, there is a need to identify which models are derived from this problematic base model. In this case, the same set of offline queries to the base model and control models can be reused across all provenance tests.
We further consider the case of reducing offline complexity in settings where offline queries cannot be reused. The current version of our provenance tester samples $T$ prompts for each parent/control model, then runs the hypothesis test to discover the most similar candidate to the tested model $g$ and shows it has significantly higher similarity. The key observation for reducing offline query complexity is that we may not need an equal number of queries to all parent/control models to identify the most similar one. If a particular parent model shows consistently higher similarity to $g$ compared to other models, we might be able to confirm it as the top candidate with fewer queries to the clearly dissimilar models. The challenge lies in determining when we have sufficient statistical evidence to conclude that one model is significantly more similar than the others, while maintaining our desired confidence levels.
This observation naturally leads us to formulate the problem as a Best Arm Identification (BAI) [5] problem in the Multi-Armed Bandit (MAB) setting. In this formulation, each parent or control model represents an “arm” of the bandit, and querying a model with a prompt corresponds to “pulling” that arm. The “reward” for each pull is the binary outcome indicating whether the model’s output matches that of the tested model $g$. The goal is to identify the arm (model) with the highest expected reward (similarity to $g$) while minimizing the total number of pulls (queries). So, we can leverage well-studied MAB algorithms that adaptively allocate queries, focusing more on promising candidates while quickly eliminating clearly dissimilar ones. The implementation of the tester based on BAI is detailed in Appendix F. Theoretical guarantees from the MAB literature could be applied to bound the number of queries needed to identify the correct parent model with high probability, but this is beyond our goals.
4 Evaluation
We evaluate our proposed provenance testing approach experimentally. Our evaluation aims to assess both the effectiveness of the approach and examine the validity of its core two assumptions. Specifically, we seek to answer the following research questions:
  1. (RQ1)
    How accurate is our provenance tester in practice and how does the number of prompts affect its performance?
  2. (RQ2)
    To what extent do derived models maintain similarity to their parents?
  3. (RQ3)
    How does the size and selection of control models impact the tester?
  4. (RQ4)
    How effective are the query reduction approaches?
Experimental Setup.
We run our model provenance testers on a Linux machine with 64-bit Ubuntu 22.04.3 LTS, 128GB RAM and 2x 24 CPU AMD EPYC 7443P @1.50GHz and 4x NVIDIA A40 GPUs with 48GB RAM. All experiments are implemented using PyTorch framework [35] and the Hugging Face Transformers library [49].
Models and Provenance Pairs.
We collect model candidates for all provenance pairs from the Hugging Face (HF) platform[21]. To avoid selection bias, we used download counts as our selection criterion, taking the most popular models subject only to hardware constraints on model size. To increase variety of candidates, we create two distinct benchmarks Bench-A and Bench-B, that differ in aspects such as model sizes, choice of pre-trained models, and ground-truth verification procedure. The full procedure of collection of models and constructions of benchmarks is described in Appendix A and their brief comparison is given in Table 1.
Table 1: Comparison of Bench-A to Bench-B on different features.
Feature Bench-A Bench-B
pre-trained models 10 57
derived models 100 383
total models 100 531
model parameters 1B-4B $<$ 1B
compilation method manual (partially) automatic
ground-truth verification higher lower
Framework.
Our default evaluation framework assumes extended model provenance testing (refer to Algorithm 2) as it allows to test for all parents at once. For sanity check, we also include an alternative framework, that considers the case where one parent is suspected (based on Algorithm 1), and this is discussed at the end of Section 4.1 and evaluated in Appendix D. We use the common value for significance parameter $\alpha=0.05$. Sampling of prompts is described in Appendix B.
Selection of control set.
In all of our provenance tests, we use the complete set of available pre-trained models from the benchmark as control models - 10 models for Bench-A and 57 for Bench-B. This selection was done to demonstrates that effective control sets can be constructed without careful manual curation or domain-specific analysis. Specifically, we make no effort to align control models with particular parent models’ domains or capabilities. We neither analyze the outputs of parent models $f$ nor attempt to match control models to specific use cases. Instead, we simply include all pre-trained models that rank among the most popular on the Hugging Face platform. This sampling approach, while simple, helps avoid introducing obvious selection bias while ensuring broad coverage of model types and capabilities. This straightforward selection strategy allows us to evaluate whether provenance testing can be effective even without carefully chosen control sets.
Figure 2: Precision and recall of the model provenance tester with different number of prompts on Bench-A (top) and Bench-B (bottom).
4.1 Accuracy of Model Provenance Tester
We evaluate the accuracy of the provenance tester by examining its performance on both Bench-A and Bench-B under different numbers of prompts. Figure 2 shows the precision and recall results from these experiments. The tester demonstrates similar performance patterns on both benchmarks, with slightly better results on Bench-A.
The precision is notably high (approximately $0.95$) when the tester uses up to $1,000$ prompts. Interestingly, however, the precision reduces as the number of prompts (test samples) increases. This is in direct contrast to common hypothesis testing, where larger sample size leads to smaller standard errors, thus higher precision. We get different results because our model provenance tester relies on detecting similarities of models. When using a smaller number of prompts, it can detect only the stronger similarities which are usually due to model provenance. However, as we increase the prompts, it starts detecting similar models that not necessarily have provenance relation. This leads to misclassification and reduced precision.
The recall behavior shows an opposite trend - it improves with a larger number of prompts, eventually reaching $80\%-90\%$ depending on the benchmark. This follows expected behavior: more prompts increase the statistical power of our hypothesis tests, enabling detection of small but significant differences in similarities. This increased sensitivity leads to higher recall rates, as the tester can detect more subtle provenance relationships that might be missed with fewer prompts.
We also examine the impact the randomness of prompt sampling on the tester’s accuracy. We conduct experiments on both benchmarks using five different randomly sampled sets of $1,000$ prompts [^3] We chose smaller number of prompts due to larger computation effort required to complete five full runs of both benchmarks. The running time is completely dominated by producing outputs from the models, which in theory is parallelizable, but in our case it was not due to limited GPU resources., with the same set of prompts used in all testers, and record the precision and recall for each run – see Table 6 of Appendix C. The results show that these values vary by $1-4\%$ between runs, indicating consistent performance across different prompt samples.
(RQ1): Our model provenance tester demonstrates high accuracy across different benchmarks, achieving precision of $90\%-95\%$ and recall of $80\%-90\%$ with $3,000$ prompts per model. Simply increasing the number of prompts does not guarantee uniformly better results, reflecting a fundamental trade-off: gains in recall might be accompanied by losses in precision.
The evaluations above are in the default framework, which assumes no candidate parent is given in each provenance test. We run similar experiments when the candidate parent is given in Appendix D. This is an easier problem (to make a wrong prediction, one needs not only to have conclusive hypothesis test that output a wrong parent, but also it should match the candidate parent), and the results confirm this: the recall of the tester in this framework is similar to the recall on the default framework, whereas the precision is very close to $100\%$.
4.2 Correctness of Assumptions
As discussed in Section 3.3, our approach relies on two key assumptions. While the high accuracy demonstrated in the previous section indirectly validates these assumptions, we provide here a detailed experimental analysis of both.
Our first assumption posits that derived models maintain significant similarity to their parent models. To evaluate this, we analyzed the similarity rankings across all provenance tests using $3,000$ prompts. For each derived model, we examined where its true parent ranked among all models in terms of similarity ratio $\mu$. The results strongly support this: in Bench-A, the true parent had the highest similarity ratio in $93\%$ of cases, while in Bench-B this occurred in $89\%$ of cases. When considering whether parents ranked in the top $50$th percentile by similarity, these percentages increased to $98\%$ and $96\%$ respectively. Thus we can conclude that our experiments indicate that derived models do indeed maintain strong similarity patterns with their parent models. Inadvertently, we have shown as well that with $3,000$ prompts the tester almost approaches the statistical limit (only the model with highest similarity ratio can be identified as a parent), as the recalls are very close to the percentages of highest similarity ($89\%$ recall vs. $93\%$ highest parent similarity, and $82\%$ recall vs. $89\%$ similarity, for the two benchmarks, respectively).
(RQ2): The assumption that derived models show significant similarity to their parent models is valid for most provenance pairs.
Our second assumption concerns whether control models can effectively establish a baseline for similarity between unrelated models. We stress that in our experiments we have chosen the control models to be simply the set of all pre-trained models in an unbiased way, without any special selection or optimization for particular parent models they are tested against. We empirically observe that such unbiased selection of control model establishes a good baseline similarity as evident from the accuracy results presented thus far.
We further examine how the size and quality of the set of control models might affect tester accuracy. We conducted experiments varying the size of the control set while keeping other parameters constant ($3,000$ prompts per test). We randomly sampled different-sized subsets from our full control sets ($10$ models for Bench-A and $57$ for Bench-B) and ran $100$ complete benchmark tests for each size, and averaged the outcomes. The results, shown in Figure 3, reveal that both size and quality of the control set significantly impact tester performance. For Bench-A, even with just $4$ control models, the tester achieved $55\%$ precision. This relatively good performance with few controls can be attributed to Bench-A consisting entirely of general-purpose, well-trained LLMs - thus any subset of these models provides a reasonable baseline for measuring similarity between unrelated models. However, for Bench-B, the randomly sampled $4$-model control set yielded less than $10\%$ precision. This poor performance stems from Bench-B containing a much more diverse set of models, including domain-specific models (e.g., for medical or coding tasks) and smaller models with varying capabilities. With such diversity, a small random subset of control models is unlikely to establish good baselines for all test cases - for instance, when testing a coding-focused model, we need coding-related models in the control set to establish proper baselines [^4] Note that in practice, unlike our random sampling experiments, one can deliberately select control models matching the domain and capabilities of the suspected parent model, thus reducing significantly the impact of size of control sets, and leaving quality of the control set as the main factor on efficiency of the tester.. Performance improves steadily as control set size increases in both benchmarks, since larger control sets are more likely to include appropriate baseline models for each test case.
Figure 3: Precision and recall of Bench-B (top) and Bench-A (bottom) with respect to smaller control set size.
(RQ3): The tester’s performance significantly degrades when the control set is too small or poorly selected.
4.3 Reducing Query Complexity
Certain pre-trained models from Bench-A and Bench-B exhibit a high degree of similarity when comparing their output tokens generated from random prompts. Table 7 of Appendix C presents the top $5$ most similar model pairs from Bench-B, measured by the percentage of matching output tokens when tested on 1,000 random prompts (column $k=1$).
To reduce the online complexity, we implement an advanced rejection prompt sampling strategy as detailed in Section 3.4. We evaluate this strategy using different parameter values $k=4,16,$ and $64$ (recall, $k$ defines how many random samples are used to produce one selected sample), comparing it to the standard provenance testing without rejection ($k=1$).
Table 7 demonstrates how the percentage of matching tokens changes with rejection sampling (columns $k=4,16,$ and $64$). For example, the most similar pair of models shows a reduction in matching output tokens from $64\%$ ($k=1$) to merely $16\%$ ($k=64$), indicating that rejection sampling significantly reduces token overlap between models. This improvement directly enhances the efficiency of provenance testing by reducing the tester’s online complexity.
Figure 4: Recall for Bench-B with different values of advanced prompt sampling defined with $k$.
Figure 4 compares the tester’s recall across different values of $k$. Notable improvements are visible even at $k=4$, with higher values of $k$ showing better results (though with diminishing returns). Specifically, the recall achieved with $1,000$ prompts at $k=1$ can be matched using only about $250$ prompts at $k=64$, representing a four-fold reduction in online complexity. Figure 5 provides a comprehensive comparison between $k=1$ and $k=64$ for both precision and recall across both benchmarks, using $4-5$ times fewer queries for $k=64$ (note, in Figure 5 the number of prompts for $k=64$ are given at the top of the plots). The results demonstrate that the tester maintains its effectiveness despite the significant reduction in queries to the tested models. For example, advanced prompt sampling achieves high levels of $90-95$% precision and $80-90$% recall while reducing the required number of prompts from $3,000$ to just $500$ per model.
Figure 5: Comparison of precision/recall for Bench-B (top) and Bench-B (bottom) when advanced online prompt sampling with $k=64$ uses four times less prompts than no advanced sampling ($k=1$).
We next evaluate strategies for reducing offline complexity, which refers to the number of queries made to pre-trained models during testing. We implement this reduction using BAI, as described in Section 3.4 and given in Algorithm 4. We test this approach on both benchmarks by setting a target budget of $T$ queries (prompts) per pre-trained model. For example, with $T=1000$ on Bench-A, which contains $10$ pre-trained models, the BAI-based provenance tester has a maximum budget of $10\cdot 1,000=10,000$ total queries to make its decision.
Table 2 compares the performance of the base tester and the BAI-enhanced version across different query budgets $T\in\{500,1000,2000\}$. The results show that the BAI tester successfully reduces offline complexity by $10\%-30\%$ (as shown in the “avg queries” column). However, this reduction comes at a significant cost to recall, while precision remains largely unchanged. For instance, with $T=1,000$ on Bench-A, BAI reduces the average number of queries from $1,000$ to $605$, but recall drops from $0.86$ to $0.63$. Similarly, on Bench-B, the average queries decrease from $1,000$ to $809$, but recall falls from $0.68$ to $0.42$. This pattern persists across different values of $T$ and both benchmarks, suggesting that the trade-off between query reduction and recall preservation is not favorable in most cases.
(RQ4): The online query optimization strategy leads to a 4-6 fold query reduction without accuracy drop, whereas the offline approach performs only marginally better and has a negative impact on recall.
Table 2: Precision and recall of the base vs BAI tester on Bench-A and Bench-B.
allowed benchmark tester avg precision recall
queries $T$ queries
500 Bench-A base 500 1.00 0.81
500 Bench-A BAI 450 0.98 0.29
500 Bench-B base 500 0.95 0.56
500 Bench-B BAI 452 0.98 0.29
1,000 Bench-A base 1,000 0.99 0.86
1,000 Bench-A BAI 605 1.00 0.63
1,000 Bench-B base 1,000 0.94 0.68
1,000 Bench-B BAI 809 0.98 0.42
2,000 Bench-A base 2,000 0.98 0.89
2,000 Bench-A BAI 1,482 0.97 0.54
2,000 Bench-B base 2,000 0.92 0.77
2,000 Bench-B BAI 1,482 0.97 0.54
5 Related Work
Model Ownership and Copyright Detection. The need for methods for determining model ownership and detecting illegitimate use of model is now well recognized; the main difference is in our formulation to tackle the concern. To determine model ownership, one type of techniques apply changes to the training dataset and model training in order to insert “watermarks” [24, 43, 53], backdoors [36, 2] or fingerprints [52, 8, 37, 28]. Specialized model ownership schemes have also been proposed for other models such as graph neural networks [55, 48]. All of these techniques, however, are orthogonal to the problem of provenance testing which we formulate in this paper. They require changes to the training of the parent model which may degrade the performance of the model. Moreover, most existing approaches do not have provable guarantees that model ownership can verified with a given confidence, so the verification is often empirically determined. Their focus is not on designing tests for determining provenance under model customizations. In particular, this is often a challenge for watermarking as they have limited robustness to typical model customizations, with some recent work acknowledging the challenge [26]. The closest work to ours is the recent work motivated by similar copyright licensing concerns [12]. The main difference to our approach is that we propose a tester with minimal assumptions of black-box access, while the prior work requires much more extra knowledge such as training and testing dataset and model parameters. Additionally, we evaluate and consider real-world customizations that are available on public repositories.
Hypothesis Tests in Machine Learning Security. Statistical hypothesis testing has been recently introduced for issues in machine learning security such as formalizing membership inference attacks via likelihood ratio tests [9, 41], auditing differentially private algorithms [30, 23], property inference [29], attribute inference [54, 17], or to proposing statistical verification methods for neural networks [6]. In this work, we also phrase the problem of provenance for fine-tuned LLMs as multiple statistical hypothesis testing, but our considered problem formulation is considerably different from those considered in prior works. There are several noteworthy differences in our formulation, compared to all of these other problem domains. In order to upper bound the hypothesis test’s false positive rate, we need to have a good estimate of the null hypothesis. For example, in some settings such as the membership inference, estimating the null hypothesis may be computationally expensive since it requires training different models under training datasets with or without given data points. Instead, in our formulation, we have multiple control models and we determine provenance depending on the reference similarity with them. As an example of another difference to privacy related inference tests, the randomness in our tests is only over the inputs given to the model, not training data points. As the choice of the input samples is independent from the training dataset of the model, our tests have soundness with respect to statistical significance.
Customization Techniques. Not many works study how fine-tuning or other customization techniques change the output distribution or the features of the pretrained model. Both training from scatch and fine-tuning optimize a similar training objective and they differ only in their initialization (random vs. pretrained weights). Because of the non-convex nature of the optimization, it is non-trivial to analytically analyze how their training dynamics converge to different minima. Some works theoretically analyze the effect of initialization (known as implicit regularization) for two-layer networks (or restricted setups), and not for pretraining [31, 42, 45]. Prior works denote that fine-tuning affects the robustness to out-of-distribution, pointing that some pretrained features get distorted in this process [27, 39, 44].
6 Conclusion
Our work formulates the model provenance testing problem for large language models which has many applications such as in detection of misuse of licensing and terms of use or problematic customized models. We present an approach based on statistical testing with minimal assumptions that has high accuracy for real-world benchmarks. Our key insight is that models derived through standard customization approaches maintain a level of similarity to their parent model that is statistically distinguishable from unrelated parents. We evaluate this observation empirically, together with our approach and several optimizations in Section 4. We find our proposed method to be practical for deciding LLM provenance.
Acknowledgments
We thank Xu Louise, Bo Wang, Jason Zhijingcheng Yu, and Mallika Prabhakar for their valuable feedback on the draft. This research was supported by the Crystal Centre at National University of Singapore.
Appendix A Models and Benchmarks
We collect model candidates for all provenance pairs from the Hugging Face (HF) platform [21]. Since there is no inherent ground truth to determine whether two models constitute a provenance pair, we employ multiple heuristic approaches. These include analyzing metadata available through the HF API and comparing the weights of downloaded models. We consider the most reliable ground truth to be cases where model uploaders explicitly specify their model as a fine-tuned version of another model, indicated by the presence of "base_model:finetune:<basemodel_name>" keyword in the model description on HF. When this explicit indication is not present, we resort to less reliable methods: we attempt to infer parent-child relationships through model naming patterns and by analyzing model descriptions on HF. Additionally, we identify potential provenance pairs by measuring the similarity between model weights, assuming that highly similar weights suggest a parent-child relationship. From these models we build two benchmarks.
The first benchmark, called Bench-A, consists of LLM pairs for model provenance constructed from popular pre-trained models and their fine-tuned derivatives. To build this benchmark, we manually selected 10 widely-used pre-trained models (refer to Tbl. 3) with between 1 billion and 4 billion parameters (the upper bound was determined by our GPU memory constraints). Among these, we purposefully included four pairs of architecturally similar models from Meta, Microsoft, Google, and AliBaba to evaluate our tester’s ability to distinguish between closely related base models and to have some control models. For each pre-trained model, we then randomly sampled 10 fine-tuned derivatives using the Hugging Face API (i.e. use highly reliable ground truth verification), prioritizing diversity in model creators. This sampling strategy resulted in 100 derived models, that constitute Bench-A.
Table 3: All 10 pre-trained LLMs from Bench-A.
Hugging Face Model # params
meta-llama/Llama-3.2-1B-Instruct 1,235,814,400
meta-llama/Llama-3.2-3B-Instruct 3,212,749,824
microsoft/Phi-3-mini-4k-instruct 3,821,079,552
microsoft/phi-2 2,779,683,840
google/gemma-2b 2,506,172,416
google/gemma-2-2b 2,614,341,888
Qwen/Qwen2-1.5B 1,543,714,304
Qwen/Qwen2.5-1.5B-Instruct 1,543,714,304
deepseek-ai/deepseek-coder-1.3b-base 1,346,471,936
TinyLlama/TinyLlama-1.1B-Chat-v1.0 1,100,048,384
Table 4: Top 10 pre-trained LLMs from Bench-B.
Hugging Face Model # params
openai-community/gpt2 124,439,808
EleutherAI/pythia-70m 70,426,624
microsoft/DialoGPT-medium 345,000,000
facebook/opt-125m 125,239,296
distilbert/distilgpt2 81,912,576
openai-community/gpt2-large 774,030,080
openai-community/gpt2-medium 354,823,168
Qwen/Qwen2-0.5B 494,032,768
JackFram/llama-68m 68,030,208
EleutherAI/gpt-neo-125m 125,198,592
The second benchmark, denoted as Bench-B, was constructed through a more automated and comprehensive approach. We began by downloading the 1,000 most popular models from Hugging Face with less than 1B parameters, ranked by download count. We then filtered out non-English models [^5] Due to lack of control models for them. and those exhibiting low entropy or high self-perplexity, which are indicators of poor training quality or insufficient learning [^6] We avoid testing low quality models.. This filtering process resulted in 608 viable models. To establish ground truth provenance relationships among these models, besides the model owners provided fine-tune keyword approach, we also used the other less reliable methods. Through this analysis, we identified 57 pre-trained models and established 383 ground-truth model provenance pairs. The remaining 148 models are considered to be independent, having no clear derivation relationship with any other models in analyzed set. Part of models from Bench-B is given in Table 4.
Appendix B Sampling Prompts
To produce prompts for our provenance testers, we use indiscriminately five popular LLMs: gemini-pro-1.5, claude-3.5-sonnet, gemini-flash-1.5, deepseek-chat, and gpt-4o-mini. Each produced prompt is an incomplete sentence containing five to twenty words – refer to Table 5 for examples.
Table 5: Examples of prompts.
1 In response to mounting public pressure, the concerned
2 The bright star known as Antares was visible even from
3 The surgeon prepared the instruments for a delicate
4 Scholars carefully examined the fragile
5 The phonetics lecturer explained the intricacies of the
Appendix C Additional Tables
Table 6: Precision and recall of the provenance tester on Bench-A and Bench-B with five different sets of 1,000 prompts.
run Bench-A Bench-B
precision recall precision recall
1 1.00 0.83 0.93 0.67
2 0.99 0.83 0.94 0.68
3 0.98 0.86 0.95 0.67
4 1.00 0.83 0.95 0.67
5 1.00 0.83 0.94 0.66
Table 7: Most similar pre-trained models from Bench-B sorted for $k=1$ (no advanced prompt sampling), and their corresponding values for $k=4,16,64$.
Model 1 Model 2 $k=1$ $k=4$ $k=16$ $k=64$
gpt2-large gpt2-medium 0.64 0.36 0.22 0.16
gpt2-large megatron-gpt2-345m 0.64 0.38 0.25 0.15
pythia-410m-deduped pythia-410m 0.62 0.37 0.24 0.15
gpt2-medium megatron-gpt2-345m 0.62 0.34 0.22 0.15
Qwen1.5-0.5B Sailor-0.5B 0.61 0.35 0.20 0.17
average 0.33 0.13 0.08 0.06
Appendix D Testing with Known Parent
We evaluate the provenance tester for cases where the parent model is given, addressing whether a pair of models $(P,C)$ constitutes a provenance pair. We construct test pairs from both Bench-A and Bench-B benchmarks. From Bench-A, we take all 100 true pairs $(P_{i},C_{i})$ and create 100 false pairs $(\tilde{P_{i}},C_{i})$ by selecting one random non-parent $\tilde{P_{i}}\neq P_{i}$ for each child $C_{i}$. This ensures a balanced dataset where random guessing would achieve 50% accuracy. We similarly obtain 766 testing pairs from Bench-B  [^7] Unlike Bench-A, Bench-B already contains (child) models that have no known parent among the 57 pre-trained models, and we use these as one of the negative pairs..
Results from testing both benchmarks are shown in Figures 6 and 7. As expected, the tester performs better when the suspected parent is known compared to cases with unspecified parents. While recall remains unchanged, precision reaches 100%. This improvement occurs because false positives now require both that the statistical hypothesis test returns the wrong parent and that this wrong parent matches the suspected parent. The precision with known parents ($P_{knownparent}$) is lower bounded by the precision with unspecified parents ($P_{unspecified}$), and depends on how we sample the incorrect parents. With uniform random sampling, it can be estimated as $P_{knownparent}=100-\frac{100-P_{unspecified}}{n-1}$, where $n$ is the number of pre-trained models. For both benchmarks, this yields precision slightly below 100%.
Figure 6: Precision and recall of running the model provenance tester on Bench-A without known parent (dark blue and dark red), and with a suspected parent (light blue and light red).
Figure 7: Precision and recall of running the model provenance tester on Bench-B without known parent (dark blue and dark red), and with a suspected parent (light blue and light red).
Appendix E Advanced Sampling of Prompts
We further give the rejection sampling of prompts used to reduce the online complexity of the tester. We present only the sampling procedure, i.e. how a set of prompts $x_{1},\ldots,x_{T^{\prime}}$ is produced given the original prompt sampler $\Omega$, and the two sets of models: the parent candidate set $F$, and the control set $C$. The tester that uses the new set is identical to the original tester, with the only difference that it samples prompts from the new set. The sampling is given in Algorithm 3.
Algorithm 3 Advanced Prompt Sampling
Parameter $k$, Candidate set $F=\{f_{1},\ldots,f_{s}\}$, set of control models $C=\{c_{1},\ldots,c_{m}\}$, prompt space $\Omega$, number of prompts $T^{\prime}$
$Prompts\leftarrow\emptyset$triangleright Initialize empty set of prompts
$H\leftarrow F\cup C$triangleright All models
$\text{same}[i][j]\leftarrow 0\text{ for all }i,j\in[|H|]$triangleright Counter of times two models produced the same output token
for $i\leftarrow 1$ to $T^{\prime}$ do
     $x_{1},\ldots,x_{k}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\Omega$triangleright Sample $k$ prompts
     $Score[j]\leftarrow 0\text{ for all }j\leftarrow 1$ to $k$triangleright Scores across $k$ prompts
     for $j\leftarrow 1$ to $k$ do $\triangleright$ Find score for each
         $s\leftarrow 0$
         for $l_{1}\leftarrow 1$ to $|H|$ do
              for $l_{2}\leftarrow 1$ to $|H|$ do
                  $old\leftarrow\frac{same[l_{1}][l_{2}]}{i-1}$
                  $new\leftarrow\frac{same[l_{1}][l_{2}]+\mathds{1}(h_{l_{1}}(x_{j})=h_{l_{2}}(x_{j}))}{i}$
                  $weight\leftarrow e^{\tau\cdot(old-new)}$
                  $s\mathrel{+}=\mathds{1}(old>new)\cdot weight$
              end for
         end for
         $\text{Score}[j]\leftarrow s$
     end for
end for
$l\leftarrow\operatorname*{arg\,max}_{j}\text{Score}[j]$triangleright Find largest score
$Prompts\leftarrow Prompts\cup\{x_{l}\}$triangleright Add that prompt
return $Prompts$
Appendix F Reducing Offline Queries with Best Arm Identification
To reduce the offline queries to the parent and control models we replace the hypothesis tests with Best Arm Identification (BAI) algorithm (that provides as well theoretical guarantees on confidence). For practical purposes, in our implementation given in Algorithm 4 we use the BAI proposed in [15].
Algorithm 4 Tester based on Best Arm Identification
Model $g$, candidate set $F=\{f_{1},\ldots,f_{s}\}$, set of control models $C=\{c_{1},\ldots,c_{m}\}$, prompt space $\Omega$, number of prompts $T$, significance parameter $\alpha$, maximum average prompts per model $N$
$M\leftarrow F\cup C$triangleright Set of all models
$U(t,\alpha):=\sqrt{\frac{\log(4t^{2}/\alpha)}{2t}}$triangleright Confidence interval for BAI
$\text{hits}[m]\leftarrow 0\text{ for all }m\in M$triangleright # same tokens with $g$
$\text{tots}[m]\leftarrow 0\text{ for all }m\in M$triangleright # queried
$A\leftarrow M$triangleright Active set of models
$t\leftarrow 0$
while True do
     $x\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\Omega$triangleright Sample prompt
     $y_{g}\leftarrow g(x)$triangleright Query $g$
     for $m\in A$ do $\triangleright$ Update hits/tots
         $y_{m}\leftarrow m(x)$triangleright query model $m$
         $hits[m]{+}=\mathds{1}(y_{g},y_{m})$triangleright Update hit counters
         $tots[m]{+}=1$triangleright Update total queries
     end for
     $\mu_{best}\leftarrow\max_{m\in M}\{\frac{\text{hits}[m]}{\text{tots}[m]}\}$triangleright Find best $\mu$
     $u\leftarrow U(t,\alpha)$triangleright Confidence radius
     for $m\in A$ do
         if $\mu_{best}-u>\frac{\text{hits}[m]}{\text{tots}[m]}+u$ then
              $A\leftarrow A\setminus\{m\}$triangleright $\mu_{m}$ too far from best
         end if
     end for
     if $|A|=1$ then $\triangleright$ Only 1 model left
         break
     end if
     if $\sum_{m\in M}\text{tots}[m]>N\cdot|M|$ then $\triangleright$ Reached max queries
         break
     end if
     $t\leftarrow t+1$
end while
if $|A|=1\text{ and }A\subseteq F$ then $\triangleright$ Model needs to be from $F$
     return $(\textsc{True},A)$
end if
return False